Opposite sides of a rectangle have the same length. If one side is represented by (2x+3) and the opposite side is represented by (5x-6) , what is the length of these sides? Please show all work.
Answer by ankor@dixie-net.com(22740)

Opposite sides of a rectangle have the same length.
If one side is represented by (2x+3) and the opposite side is represented by (5x-6) , what is the length of these sides?
therefore
5x - 6 = 2x + 3
5x - 2x = 3 + 6
3x = 9
x = 3
then one side is:
5(3) - 6 = 9
and you can find the other side